# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

### v1

@@ 7. Conclusion and Future Work
In conclusion, this study introduces bank statement transaction data as alternative data for MSMEs loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature of alternative data. In the future, we aim to integrate real-time transactional data into our workflow to better assess the creditworthiness of MSMEs with limited credit history in Malaysia.

### v2

@@ 8. Conclusion
In conclusion, this study introduces bank statement transactions as alternative data for MSMEs loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature and practical value of alternative data.
Acknowledgements.
We thank the Cradle Fund and the Microsoft for Startups program for their continuous support. This work was made possible through the contributions of the DocLab and DocSuite teams at AI Lens.